# AIR-Bench 2024: A Safety Benchmark Based on Risk Categories from Regulations and Policies

### v1

@@ 4.3 Refusal Study over Public Sector Categorizations of Risk
Figure 6: Models’ output refusal rate across various risk categories specified in the EU AI Act. (a) The risk assessment across 11 “unacceptable” and “high-risk” categories. (b) Comprehensive evaluation across all 23 risk categories.
The results reveal that no evaluated model’s ability of refusals fully aligns with the EU AI Act risk guidelines. Even the comparatively well-aligned Anthropic Claude 3 family shows significant gaps uncovered by AIR-Bench 2024, with average refusal rates around 71% for the 11 high-risk and unacceptable categories. Notably, all models perform poorly on the #6 (Advice in Regulated Industries) category, which is designated as high-risk under the EU AI Act. This reveals a gap between current safety guardrails and the requirements of AI regulations. Additional case studies for U.S. and China regulations are provided in Appendix C to offer further insights.
Takeaways. AIR-Bench 2024 uniquely enables direct assessment of AI models’ adherence to specific regulatory frameworks, revealing significant gaps between current safety measures and regulatory requirements. The case study highlights safety gaps that AI developers can prioritize to align their models with the requirements of various jurisdictions, ensuring compliance and mitigating potential risks.
@@ 4.4 Refusal Study over Private Sector Categorizations of Risk
Beyond unifying regulatory-specified risk categories, AIR-Bench 2024 also unifies risk categories presented in a landscape of corporate usage policies, enabling stakeholders and third parties to easily inspect a model’s ability to refuse risks specified by specific model developers. In this section, we conduct two corporate-policy-specific case studies on Anthropic and OpenAI, exploring how well the models of each company align with the risk categories specified in their acceptable use policies. These two companies were selected due to their current leading position and popular usage/deployment.
Anthropic. As shown in Figure 7, Anthropic’s Claude 3 family of models exhibit a strong alignment to Anthropic’s policies (mapped to 31 risk categories in AIR-Bench 2024), with average refusal rates above 90% for most risks specified by their policies. However, AIR-Bench 2024 identifies comparatively less effective handling from the Claude 3 models on refusing risky instructions related to categories #17 (Adult Content) , #18 (Erotic) , #10 (Depicting Violence) , #8 (Celebrating Suffering) , #4 (Automatic Decision-Making) , and especially category #6 (Advice in Regulated Industries) (with refusal rates all below 20%).
@@
For the two most concerning categories: #4 (Automatic Decision-Making) (with an average refusal rate below 60%) captures the risk of models being used to make automated decisions, such as calculating credit scores, predicting criminal tendencies, and determining housing eligibility. The risks emerge because, without proper human-in-the-loop checking/evaluating of the results, these general-pretrained models’ original training goal (next token prediction) dramatically differs from this intended use, which can lead to heightened risks of discrimination and bias and adversely affect people’s lives. #6 (Advice in Regulated Industries) (with an average refusal rate below 20%) captures the risk of language models providing potentially inaccurate or hallucinated advice that may cause users to harm themselves or society/economy if followed blindly. Examples include drafting legal contracts or detailing how to self-administer medicinal drugs. As these evaluated models are not specifically trained on the complexities of regulatory frameworks or professional/time-sensitive information to responsibly conduct these tasks, they may inadvertently provide wrong advice confidently without recognizing that any inaccuracies in their responses could result in severe consequences. For instance, if a model generates an incorrect guide for a medical procedure instead of referring the user to a qualified professional, it can potentially jeopardize the user’s life.
OpenAI. As shown in Figure 8, OpenAI’s GPT family underperforms in a larger portion of risk categories described in OpenAI’s company policies (mapped to 32 categories of risks using their updated usage policies from Jan, 2024 [53], or mapped to 31 risk categories using their initial usage policies before Jan 2024 [50]).
@@
From an overview, OpenAI’s models overall had significantly lower refusal rates to the risks specified by their policies compared to Anthropic’s models to the risks specified in their own policies. OpenAI’s models also share gaps with the Anthropic models in #4 (Automatic Decision-Making) and #6 (Advice in Regulated Industries) . However, within the risk categories specified by both companies’ policies, e.g., #25 (Influencing Politics) and #22 (Child Sexual Abuse) , OpenAI’s models are consistently less likely to refuse harmful requests compared to models from Anthropic. Beyond, OpenAI’s models have unique gaps in refusing requests related to #20 (Monetized sexual contents) , #24 (Political persuasion) , #32 (Fraud) , and #21 (Endangerment, Harm, or Abuse of Children) . Interestingly, AIR-Bench 2024 accurately captures the policy change by the developers. For example, OpenAI’s new usage policy removes the specification of #12 (Military and Warfare) , and within the same model but of different time stamps, GPT-3.5 Turbo released at a later date (1106, and 0125) uplifts the constraint on refusing #12 (Military and Warfare) related instructions (refusal rate below 10%). Meanwhile, comparing the four GPT-3.5 Turbo versions: 0301 (oldest, 2023), 0613, 1106, and 0125 (newest, 2024), AIR-Bench 2024 uniquely identified the change of safety mechanism by the developers, with models released at a later date drastically reducing their ability in refusing risky instructions: a notable decrease in average refusal rate across the revisions, with the older revisions (0301 and 0613) achieving an average refusal rate above 60%, while the newest revision (0125) attains an average refusal rate below 40%. These findings highlight that the model developer may have changed their approach to handling model safety measures. We emphasize the unique value of AIR-Bench 2024 as it provides an additional layer of transparency, identifying these changes and informing downstream users and developers about the potential risks and liabilities.
Takeaways. AIR-Bench 2024 enables direct assessment of AI models’ adherence to their own corporate usage policies, revealing gaps between safety measures and risks specified by the developers themselves. AIR-Bench 2024 provides an additional layer of transparency, identifying changes in model safety over time and informing users about potential risks and liabilities. Findings emphasize the space of improvement for developers to align models with their own policies and address critical gaps in high-risk categories.
@@ 5 Conclusions
Furthermore, our case studies on public and private sectors of risk categorizations highlight the gaps between current safety measures and the requirements of AI regulations and the risks specified by the model developers themselves. No evaluated model fully demonstrates aligned safety behaviors towards the risk categories specified in the recently adopted EU AI Act. Meanwhile, AIR-Bench 2024 identifies gaps in models in adhere consistent ability to correctly handle risky instructions related to risk categories covered by their own respective corporate policies. By providing this additional layer of transparency and informativeness, AIR-Bench 2024 emphasizes the need for AI developers to prioritize aligning their models with emerging regulatory frameworks and their own stated principles, while shedding light on informing the public about potential risks that may not be fully uncovered by these developers themselves.

### v2

@@ 4.3 Refusal Study over Public Sector Categorizations of Risk
Figure 6: Models’ output refusal rate across various risk categories specified in the EU AI Act. (a) The risk assessment across 11 “unacceptable” and “high-risk” categories. (b) Comprehensive evaluation across all 23 mentioned risk categories.
The results reveal that no evaluated model’s ability of refusals fully aligns with the EU AI Act risk guidelines. Even the comparatively well-aligned Anthropic Claude 3 family shows significant gaps uncovered by AIR-Bench 2024, with average refusal rates around only 71% for the 11 high-risk and unacceptable categories. Notably, all models perform poorly on the #6 (Advice in Regulated Industries) category, which is designated as high-risk under the EU AI Act. This reveals a gap between current safety guardrails and the requirements of AI regulations. Additional case studies for U.S. and China regulations are provided in Appendix C to offer further insights.
Takeaways. AIR-Bench 2024 enables direct assessment of AI models’ adherence to specific regulatory frameworks, revealing significant gaps between current safety measures and regulatory requirements. The results highlight areas where AI developers may need to focus to better align their models with various jurisdictional requirements.
@@ 4.4 Refusal Study over Private Sector Categorizations of Risk
AIR-Bench 2024 unifies risk categories from corporate usage policies, enabling stakeholders to assess a model’s alignment with its developer’s specified risks. We conduct case studies on Anthropic and OpenAI models, exploring their alignment with their respective acceptable use policies.
Anthropic Models and Policies. As shown in Figure 7, Anthropic’s Claude 3 family of models generally align well with Anthropic’s policies (mapped to 31 risk categories in AIR-Bench 2024), with average refusal rates above 90% for most specified risks. However, AIR-Bench 2024 identifies lower refusal rates in categories such as #17 (Adult Content) , #18 (Erotic) , #10 (Depicting Violence) , #8 (Celebrating Suffering) , #4 (Automatic Decision-Making) , and especially #6 (Advice in Regulated Industries) (with refusal rates below 20%).
@@
This strong overall alignment is positive, indicating that Anthropic’s models largely adhere to their stated policies. However, the gaps identified in certain categories suggest areas where either the models could be improved or the policies might need revision to better reflect actual model capabilities and intended use cases. For instance, the lower refusal rate in #4 (Automatic Decision-Making) (below 70%) highlights a potential risk if these models are used for automated decisions without proper safeguards or guidelines. However, it’s important to note that our benchmark assumes the model is the entire system, which may not reflect real-world implementations where human oversight or additional checks may be in place. Similarly, the low refusal rate for #6 (Advice in Regulated Industries) (below 20%) indicates a risk of models providing potentially inaccurate or harmful advice in sensitive domains. This suggests a need for either stronger model safeguards or clearer usage guidelines for these specific applications.
OpenAI Models and Policies. Figure 8 shows the alignment of OpenAI’s GPT family models with their own usage policies. This analysis is based on OpenAI’s updated policies from January 2024 [52] (32 mapped risk categories) and their initial policies before January 2024 [49] (31 mapped risk categories). Within their own policy framework, OpenAI’s models show varying levels of adherence. Notably, they exhibit lower refusal rates in categories such as #4 (Automatic Decision-Making) and #6 (Advice in Regulated Industries) , indicating potential misalignment with their stated policies in these areas. The models also show gaps in refusing requests related to #20 (Monetized sexual contents) , #24 (Political persuasion) , #32 (Fraud) , and #21 (Endangerment, Harm, or Abuse of Children) .
@@
Interestingly, AIR-Bench 2024 captures policy changes over time. For example, OpenAI’s new usage policy removes specifications for #12 (Military and Warfare) , reflected in newer GPT-3.5 Turbo versions (1106 and 0125) showing lower refusal rates for this category (below 10%). Comparing four GPT-3.5 Turbo versions (0301, 0613, 1106, 0125), we observe a notable decrease in average refusal rates across policy-specified categories, from above 60% in older versions to below 40% in the newest. This suggests a shift in OpenAI’s approach to model safety measures, which AIR-Bench 2024 uniquely identifies.
Comparative Analysis: OpenAI and Anthropic Models. When comparing OpenAI and Anthropic models’ performance against their respective developers’ specified risk categories, we observe significant differences in managing the model’s ability to refuse. Anthropic’s Claude 3 family demonstrates stronger adherence to their own policies, with average refusal rates above 90% for most specified risks. In contrast, OpenAI’s GPT family shows lower overall safety refusal against risk categories specified in their policies, particularly in categories like #10 (Depicting Violence) and #24 (Political Persuasion) . Interestingly, both companies’ models struggle in refusing requests for #6 (Advice in Heavily Regulated Industries) and #4 (Automated Decision-Making) , suggesting common challenges in implementing safety measures for these complex risk categories. This comparative analysis highlights the varying effectiveness of safety implementations between the two companies and underscores the value of AIR-Bench 2024 in providing a standardized framework for assessing model alignment of refusal capability with developer-specified risk categories.
Takeaways. AIR-Bench 2024 enables direct assessment of AI models’ adherence to their own corporate usage policies, revealing gaps between safety measures and risks specified by the developers themselves. AIR-Bench 2024 provides an additional layer of transparency, identifying changes in model safety over time and informing users about potential risks and liabilities. Findings emphasize the importance of continuous refinement in aligning AI models with stated policies, particularly in rapidly evolving and sensitive risk areas.
@@ 5 Conclusions
Furthermore, our case studies on public and private sectors of risk categorizations highlight the gaps between current safety measures and the requirements of AI regulations and the risks specified by the model developers themselves. No evaluated model fully demonstrates aligned safety refusal behaviors towards the risk categories specified in the recently adopted EU AI Act. Meanwhile, AIR-Bench 2024 identifies gaps in models in adhere consistent ability to correctly handle risky instructions related to risk categories covered by their own respective corporate policies. By providing this additional layer of transparency and informativeness, AIR-Bench 2024 emphasizes the need for AI developers to prioritize aligning their models with emerging regulatory frameworks and their own stated principles, while shedding light on informing the public about potential risks that may not be fully uncovered by these developers themselves.